### v2

# Clocking and controlling attosecond currents in a scanning tunnelling microscope

Clocking and controlling attosecond currents in a scanning tunnelling microscope
Quantum tunnelling of electrons can be confined to the sub-cycle time scale of strong light fields 1, 2, contributing decisively to the extreme time resolution of attosecond science 3. Because tunnelling also enables atomic-scale spatial resolution in scanning tunnelling microscopy (STM 4), integrating STM with light pulses has long been a key objective in ultrafast microscopy, spanning the picosecond 5, 6, 7, 8 and femtosecond domains 9, 10, 11, with first signatures of attosecond dynamics 12. However, while sub-cycle dynamics on the attosecond time scale are routinely controlled and determined with high precision 1, 13, 2, 14, 15, 16, 17, 18, 19, controlling the direction of attosecond currents and determining their duration have remained elusive in STM. Here, we induce STM tunnelling currents using two-colour laser pulses and dynamically control their direction, relying solely on the sub-cycle waveform of the pulses. Projecting our measurement data onto numerical and analytical solutions of the time-dependent Schrödinger equation reveals non-adiabatic tunnelling as the underlying physical mechanism, yielding a current burst duration of 860 as. Despite working under ambient conditions but free of thermal artifacts, we achieve sub-angström topographic sensitivity and a lateral spatial resolution of 2 nm. This unprecedented capability to directionally control attosecond bursts will enable triggering and imaging ultrafast charge dynamics in atomic, molecular and condensed systems at the spatio-temporal microscopy frontier of lightwave electronics 20, 21, 22.
Introduction
Attosecond science is based on the control of electron motion on the atomic time scale by strong electric fields 3. While it has enabled the observation of a wide range of attosecond phenomena, achieving a spatial resolution on the order of the atomic length scale in the angström-nanometer regime is challenging. The main pathway to attosecond microscopy is the use of attosecond-scale electron pulses instead of light due to their small de Broglie wavelength, typically much less than 1 nm. In addition, electrons must be spatially confined, either as a recolliding electron wavepacket originating from an atom, molecule, or nanostructure 23, 24, 19, 18, as a narrow beam in an electron microscope 25, 26, or inside an atomic-scale tunnelling microscopy (STM) junction 12, 27, 11. The strength of the latter approach is that STM in its static implementation using a bias field provides atomic resolution and precise energy resolution out of the box for a wide range of systems, such as molecules 28. STM tunnelling currents driven by single-cycle THz transients 5, which can be regarded as a slowly varying static bias field, have enabled real-time observations of atomic-scale molecular vibrations on the sub-picosecond time scale 6 and exciton formation in molecules 8, for instance.
Tuning the carrier-envelope phase (CEP) can shape an ultrashort THz field waveform to break its symmetry and obtain a field that is much stronger in one direction compared to the other. Maximum tunnelling is then achieved by the field in that direction, enabling control of the dominant direction of the current. This allows, for example, the controlled extraction of an electron from the sample 7, an important prerequisite for pump-probe measurements 6. Moving to infrared wavelengths and the femtosecond and attosecond time scales, CEP modulation is a well-known approach to control the sub-cycle electron motion in atoms 29, 30, 1, metallic nanotips 13, 31, 18, nanodevices 17 and vacuum nanogaps 15, 16, 32 and enables attosecond time resolution in these systems. It may seem straightforward to induce attosecond tunnelling currents and control their direction in STM using CEP-stable infrared laser pulses, building on many experiments exploring ultrafast STM in the femtosecond regime (see, e.g., 9, 33, 10, 34, 35, 36, 11). However, pioneering ultrafast STM experiments with infrared lasers have shown that thermal artifacts are often present when conventional lock-in approaches are used to extract low-level laser-induced currents 9, 37. First signatures of CEP-controlled tunnelling current modulation in STM using 6-fs near-infrared laser pulses were reported recently 12. Still, sub-cycle control of the direction of STM currents and timing of their dynamics have remained elusive. This advance is necessary for attosecond imaging of ultrafast electron phenomena.
In this work, we demonstrate robust attosecond directional control of ultrafast tunnelling currents in an STM junction. We employ two-colour pulses, a powerful alternative to CEP-stable pulses, and achieve ultrafast measurements without thermal artifacts. To this end, we superimpose an infrared laser pulse with its second harmonic in the same polarization plane and change their relative time delay. Two-colour pulses lead to field waveform with controlled symmetry breaking and enables sub-cycle control, as demonstrated in a wide range of attosecond phenomena 38, 39, 40, 41, 19. Using the synthesized field waveform of the two-colour pulse, we are able to seamlessly switch between tip-to-sample tunnelling and vice versa. From an excellent agreement of experiment and theory we are able to identify the underlying physical mechanism as non-adiabatic tunnelling, with electron burst durations of as. Free of thermal artifacts, our microscope enables a lateral spatial resolution of 2 nm despite operation under ambient conditions.
Results
Ultrafast two-colour scanning tunnelling microscopy
In our experiment, we perform STM at ambient conditions at room temperature using two-colour laser pulses (see Fig. 1a). An electropolished Pt:Ir nanotip with an apex radius of 5…20 nm and a gold substrate form a tunnelling junction with variable width . We irradiate the junction with 35 fs, 1850 nm laser pulses of up to 75 pJ pulse energy at a repetition rate of 80 MHz together with its second harmonic (SH). The intensity ratio of the SH and the fundamental is 10% in order to induce a pronounced symmetry breaking of the resulting waveform (see the inset of Fig. 1a for an illustration). Figure 1b shows the optical setup of our experiment (see Methods for a detailed description). We control the relative time delay between the two colours using an interferometric setup with the help of a linear delay stage. The two-colour pulses induce an optical near-field in the nano-scale tunnelling junction, leading to a strong field enhancement and enabling us to enter the strong-field sub-cycle regime of light-matter interaction.
The microscope is operated in constant-current measurement mode, where a static bias field drives a tunnelling current and a feedback loop controls the relative distance between tip and sample to achieve a constant current. In addition, ultrafast laser-induced currents are generated by the optical near-field formed inside the tunnelling junction through irradiation with the two-colour laser pulses. Isolating the contribution of the laser-induced currents from the static tunnelling current is a challenging task. A chopper-based lock-in approach (cf. 12) has not been applicable because it produces artifact signals. Periodic on-off switching of the intensity in the kHz domain leads to expansion and contraction of the nanotip and a corresponding increase and decrease of the static tunnelling current, masking the actual laser-induced current signal (see Supplementary Information). In order to avoid any intensity modulation and the artifacts accompanying it, we modulate the two-colour delay with a sinusoidal modulation 34 at an angular frequency of kHz using a mirror mounted on a vibrating piezoelectric chip. The total two-colour delay as a function of time is then given by , where is the base delay set by the linear delay stage and is the amplitude of the sinusoidal delay modulation. This leads to a clear and consistent lock-in signal and allows us to measure the laser-induced current independently of the microscope feedback loop operating at a locking bandwidth below 500 Hz.
Figure 1c-e shows a scan of the gold sample under laser irradiation for a fixed value of . The sample bias voltage is 200 mV and the set-point current for the conventional STM operation is 100 pA. The topography map (Fig. 1a) obtained from the static tunnelling current reveals fine atomic steps of the gold sample on the angström scale. Simultaneously with the topography, the lock-in signal at 3.7 kHz allows us to record a map of the laser-induced current (Fig. 1d), which largely follows the topographic features of the sample. The phase offset of the laser-induced current signal with respect to the delay modulation is found to be approximately uniform (see Supplementary Information). A look at a line cross-section (see Fig. 1d) shows that the laser-induced current signal is sensitive to the tiny variations of the topographic height of the sample on the sub-angström scale. We find that peaks in the signal are correlated with small topographic protrusions, hinting at a remarkable sensitivity of the optical near-field to the atomic-scale sample geometry at the STM junction 42, 43. We estimate the lateral spatial resolution obtained from the laser-induced signal to be around 2 nm, roughly the same as the spatial resolution of the conventional topographic scan. Here we are likely limited by the fact that the microscope is operated in ambient conditions.
Sub-cycle waveform control of tunnelling currents
Scanning the base two-colour delay with a step size of 130 as and placing the nanotip at a fixed spot on the sample allows us to study the sub-cycle dependence of the laser-induced current. Figure 2a shows the lock-in current as a function of (see Supplementary Information for a plot of a wide-range delay scan). We find oscillations with a period of 1.5 fs, which corresponds to half of the period of the SH field, a clear sub-cycle feature. The lock-in measurement also yields the phase of the lock-in current signal (Fig. 2b). We find alternating phase jumps of 180 degrees on top of an arbitrary but constant phase offset. A procedure to reconstruct the actual laser-induced current from the lock-in signal (see Supplementary Information for details) reveals that controls the magnitude of the current, as shown by the dashed orange curve in Fig. 2c; it may also indicate control of the direction of the current. However, the reconstruction from the lock-in measurements alone cannot provide any information about the offset of the current which remains undefined. A different, more direct measurement of the laser-induced current is required.
In order to show that indeed controls the direction of the laser-induced current, we temporarily switch off the microscope feedback loop as well as the lock-in modulation of the two-colour delay (). We freeze the position of the nanotip in space and set the sample bias to zero, thus suppressing the static tunnelling current and preempting any symmetry breaking by the static field. The only measurement quantity that remains is the net laser-induced current, now measured directly. The red curve in Fig. 2c shows the resulting signal, which we find to be oscillatory with the period of the SH field. The signal is in good agreement with the laser-induced current reconstructed from the lock-in approach. Our measurements unambiguously demonstrate that the direction of the current is governed by . Positive currents indicate that the asymmetric waveform of the two-colour field drives electrons in sub-cycle bursts from the nanotip to the sample. For negative currents, the opposite is true – we obtain a current from the sample to the nanotip. Here, it is solely the laser field that breaks the symmetry of the tunnelling junction and produces a net current, a hallmark of lightwave electronics, where the sub-cycle waveform of a laser pulse governs ultrafast electric current dynamics 21. Since our laser pulse comprises many cycles, we obtain a train of attosecond current bursts with a periodicity of the fundamental field.
Figure 2d shows the polarization dependence of the laser-induced current measured at zero bias while the nanotip is frozen. The polarization angle is defined with respect to the sample plane. The laser-induced current peaks sharply at a polarization angle of 100 degrees, which coincides approximately with the scenario in which the two-colour laser field is aligned with the perpendicularly oriented nanotip. This ensures maximum near-field enhancement in the gap and therefore maximum laser-induced current. At 280 degrees, the same scenario is reached, but the flip of the polarization causes a reversal of the sign of the two-colour waveform. For the two angles, we perform a two-colour delay scan and indeed observe the resulting reversal of the direction of the current (Fig. 2e). This further corroborates the notion that the waveform controls the transport of electrons through the junction.
Electron transport in the non-adiabatic tunnelling regime
A crucial question which needs to be answered is the underlying physics of the electron transport. An ultrafast STM experiment 12 and theory studies 44, 45, 46 show that the dichotomy of a multiphoton regime and a laser-induced tunnelling regime, parametrized by the Keldysh parameter 47 , also governs ultrafast STM. In STM, the only measurement observable is the average current, unlike in nanotip photoemission experiments, for instance, where the photoelectron spectrum can reveal crucial information about the underlying physics (see, e.g. 13, 31, 19). Here we measure the scaling of the laser-induced current with total laser power (SH and fundamental), shown in Fig. 3a in a double-logarithmic scale. Freezing the tip at zero bias, we determine the power scaling of the amplitude of the two-colour current modulation for three different junction widths, free from any other signals. We observe that the current increases approximately like a power law with a slope of , which may indicate the absorption of 3 or 4 photons from the two-colour pulse. However, beyond a power of 4 mW, the nonlinearity begins to decrease. A soft kink appears in the scaling, a signature of the transition to the tunnelling regime of ultrafast STM 12, 44, analogous to photoemission from nanostructures 48, 49. Another effect that may contribute to the kink is the fact that we begin to transfer more than one electron per laser pulse starting at an average current of 13 pA.
In order to gain a deeper insight into the physical mechanism, we employ a simple theory model, a numerical integration of the time-dependent Schrödinger equation (nTDSE, see Supplementary Information for details) and directly compare the experiment with our calculations, taking into account all known experimental parameters (see Fig. 4a). Here, we account for the fact that the optical near-field enhancement, which translates the total laser power into the actual field driving electrons across the tunnelling junction, decreases with increasing junction width. The calculated curves allow us to extract field enhancement factors for the fundamental field of , and for 1 nm, 1.5 nm and 2 nm, respectively. These values are qualitatively in agreement with numerical solutions of the Maxwell equations for the tip-sample geometry (see Supplementary Information). The excellent agreement between experiment and simple single-electron theory allows us to determine the effective Keldysh parameter (see the Supplementary Information for the derivation of for a two-colour field). The upper horizontal axis in Fig. 3a shows for a junction width of 1 nm. We find that the dominating mechanism is non-adiabatic tunnelling in the vicinity of . Here, an electron tunnels through the potential barrier strongly suppressed by the peak of the near-field (see Fig. 3b for an illustration for electron transport from the tip to the sample). However, due to the fact that the barrier cannot be considered as quasi-static in this regime and varies strongly during the evolution of the tunnelling process, the electron is gaining energy “under the tunnelling barrier” by photon absorption (cf. 50, 51, 46). In our case, non-adiabatic tunnelling is the first step of the three-step mechanism of ultrafast STM 44, 45, 46. In the second step, the electron exits the tunnelling barrier, is further accelerated, and is transmitted into the adjacent metal in the third and final step.
Timing the attosecond current bursts
The sub-cycle waveform control of the direction of the attosecond current bursts can also be explained by the nTDSE model and by an analytical strong-field (SF) model based on the van Vleck propagator 46 (see Methods and Supplementary Information for details). Figure 4a shows the laser-induced current measured from Fig. 2c. In addition, we display the theory curves obtained from the nTDSE and SF models for a peak field strength of of the fundamental () after scaling their amplitude to match the experimental data. We find excellent agreement. The experimentally observed current modulation is nearly perfectly symmetric with respect to zero current, which attests to the fact that the symmetry breaking induced by the two-colour waveform dominates over all other symmetry-breaking factors, for example the stark difference in geometry between nanotip and sample.
The excellent overall agreement between experiment and theory allows us to trace the electron dynamics and estimate the duration of the attosecond electron bursts. Figure 4b shows the current density map calculated with the nTDSE as a function of space and time across a 1 nm junction for a peak field strength of () and for maximum symmetry breaking. The electrons are created at the peak of the field (time zero) and cross the barrier on a scale much shorter than the optical cycle duration and hence without any quiver motion 16, 44, 45. The flow of the current density matches well with semiclassical trajectories obtained from the SF model that appear after the tunnelling exit located at 0.35 nm (see dotted curves in Fig. 4b). The full-width at half maximum (FWHM) duration of the electron burst at the end of the junction at nm is depicted in Fig. 4c and amounts to 860 as. This is shorter than 970 as obtained in the single-colour case (1850 nm field only), showing that the two-colour approach not only induces full symmetry breaking, but also leads to a shorter burst duration within the sub-femtosecond time domain. In Fig. 4c, we can also observe that the arrival time is about 700 as after the peak of the field, which accounts for the travel time through the junction. The burst duration is strongly confined to the sub-cycle regime for a wider range of peak field strengths (see Fig. 4d), showing that attosecond current bursts can be generated routinely. The same is true for variations of more experimental parameters, such as the intensity ratio of the SH and the fundamental, junction width and tip workfunction; the burst duration stays well below the 1 fs mark (see Supplementary Information). While our experiment is performed at an intensity ratio of the SH and the fundamental of 10%, ratios as small as 2.5% can induce strong symmetry breaking (see Supplementary Information).
Conclusion
Our work shows that two-colour sub-cycle waveform control can induce a train of sub-femtosecond current bursts across a variable STM junction and robustly control their direction, allowing for a lateral spatial resolution of 2 nm and the ability to sense sub-angström topographic features under ambient conditions and free from thermal artifacts. In the future, our approach can be combined with all-optical readout 43 to directly reveal the evolution of the sub-cycle currents. Moving from a two-colour laser pulse to a single-cycle infrared pulse will naturally lead to a single isolated attosecond electron burst with waveform-controlled timing and direction. Combining the ability to generate these bursts with the well-established capability to image electronic states in molecules, defects, and nanostructures using conventional STM in ultrahigh vacuum conditions, we envision the combined real-time and real-space observations of coherent electron-hole dynamics and many-body effects as they unfold. For example, a sub-cycle controlled infrared pump pulse will enable extracting an electron from a molecule with attosecond precision in time at a specific atomic site, inducing a coherent evolution of complex electron-hole dynamics. A second, sub-cycle controlled probe pulse with opposite field waveform can then inject an electron back into the molecule, allowing for a readout of the dynamics before dephasing effects set in. With attosecond STM, microscopy has reached its ultimate spatiotemporal resolution limits at the atomic scale.
Methods
Experimental setup
100 fs laser pulses from an Er:fiber laser system (Menlo C-Fiber High Power, 500 mW, 1550 nm, 80 MHz) are focused into a fiber assembly (standard polarization-maintaining fiber of 75 mm length spliced to 12 mm of highly nonlinear fiber, Thorlabs HN1550P) for supercontinuum generation spanning from 950 nm to 2100 nm (cf. 52). We use only the solitonic part of the resulting supercontinuum with a central wavelength of 1850 nm by rejecting the spectral components below 1500 nm using a dichroic filter. These pulses are focused into a bismuth triborate (BiBO) crystal with a thickness of 1 mm for second harmonic (SH) generation, avoiding the messy part of the spectrum below 1650 nm by an appropriate choice of phase-matching conditions. The recollimated fundamental and SH beams are then fed into an interferometric setup where they are separated by a dispersion-engineered dichroic mirror. The SH beam passes through several filters to remove any remaining trace of the fundamental beam. Subsequently, its polarization is rotated to match the polarization of the fundamental. It then hits a mirror glued onto a piezo chip at a near-zero angle of incidence. The piezo can be modulated at kHz frequencies using a high-voltage signal to periodically change the temporal delay of the SH beam, providing the modulation to produce the lock-in signal. The delay modulation amplitude in our experiment is fs. The fundamental beam is passed through 1 mm of silicon to compensate for the dispersion of the other optical elements in the system and a longpass filter at 1650 nm to remove the messy part of the spectrum below that wavelength. We also control the base two-colour delay using a mirror mounted on a linear closed-loop piezo stage at a near-zero angle of incidence. Both beams are recombined by another dispersion-engineered dichroic mirror and expanded by a factor of 3. We monitor the two-colour delay using an auxiliary collinear interferometer setup and correct for drifts.
Before sending the beams to the microscope (RHK Technologies PanScan Flow Kit operated in ambient conditions), we control the laser power of the combined beams using a variable neutral-density filter and tune their joint polarization by an achromatic half-wave plate. An off-axis parabolic mirror with a focal length of 50 mm focuses the beams on the tip-sample junction at an angle of incidence of 20 degrees with respect to the sample plane. We use commercial Pt:Ir nanotips (Unisoku P-100PtIr(S)) with apex curvature radii of 5…20 nm and a flat gold substrate fabricated according to a self-formation procedure 53. The fundamental laser beam is focused to a spot size of 17 m. The duration of the pulse is 35 fs. The SH beam is focused to a spot size of 10 m. The duration of the SH pulse is 62 fs due to uncompensated chirp. Taking into account all parameters, the ratio of the SH intensity to the fundamental intensity is about 10 %. The gold sample can be biased, and the current is measured using a current preamplifier connected to the tip. In the tip freezing experiments, we switch off the bias and the microscope feedback loop for a duration of 600 ms and take several data points during this time.
Theory modeling
Our theory modeling is based on the one-dimensional time-dependent Schrödinger equation (TDSE) in the single active electron approximation. Despite the strong degree of simplification, such an approach has yielded excellent agreement with nanotip experiments 13, 19, for instance. Tip and sample are modeled as potential wells and populated with a wavefunction at the Fermi level (see Supplementary Information for more details). We follow two approaches, a numerical integration of the TDSE (nTDSE) based on the Crank-Nicolson method (see Supplementary Information) and an analytical strong-field (SF) model. In the following, we briefly describe the SF model 44, 46.
Based on the Dyson equation, we define the tunnelling amplitude from tip to sample as
where and are time-dependent wavefunctions in the junction and the sample, respectively. The notation at the end of Eq. 1 indicates the subtraction of the term inside the brackets evaluated at from the term evaluated at .
In analogy to the strong-field approximation of attosecond science, the wavefunctions and can be obtained by using the Van Vleck propagator and the eigenfunction of the sample system 46. The simplified tunnelling amplitude is now
Here, is a prefactor originating from the transition matrix elements. For an initial bound state in the tip and a final state in the sample, the action in the exponent is given by
where is the effective canonical momentum, is the vector potential of the laser field, is the junction width, and is the image potential. We use and to represent the emission time and the arrival time of the electron transport from tip to sample.
Following the saddle-point technique with , we can obtain the three saddle-point equations:
Here, the subscript (s) indicates that these values correspond to complex-valued mathematical saddle points rather than physical quantities. is the image potential averaged over the length of the junction. The three equations above provide a three-step framework for describing electron transport. (1) At the moment of emission, the bound electron is released from the tip by the laser field. This process is governed by the energy condition in Eq. 4. (2) After emission, the electron is accelerated by the laser field as it travels from the tip toward the sample. Its motion is described by Eq. 5. Unlike the characteristic recollision dynamics in attosecond science, the electron does not return to its origin. (3) Transmission into the sample: The energy accumulated during the second step is transferred to the sample, as described by Eq. 6. In Newtonian mechanics, the displacement of a classical point-like electron in the laser field is
Since the saddle points are complex, the integral must be taken also over a complex contour. In order to display the resulting semiclassical trajectories in Fig. 4b, we take the real part of .
Data availability
The data that support the plots within this paper and other findings of this study are available from the authors upon request.
Code availability
The codes that support the findings of this study are available from the authors upon request.
Author contributions
D. D., B. M. and A. G. contributed equally to this work.
M. K. conceived and supervised the project. A. G., S. C., B. M. and M. K. designed and built the experimental setup. D. D., A. G., B. M. and M. K. performed the measurements. B. M. carried out the theory calculations. M. K. wrote the initial manuscript. All authors contributed to the preparation of the final manuscript.
Corresponding authors
Correspondence to Michael Krüger.
Acknowledgements
The authors thank A. Borisov, M. Ivanov, A. Kollin and Y. Bekenstein for insightful discussions, I. Kaminer for providing lab space for initial experiments, A. Feigenboim for supporting the design of experimental system, I. Sthzeglowski for the initial setup of the fiber laser, P. Sidorenko for assistance with fiber splicing, and N. Dudovich, U. Leonhardt and L. M. Procopio for providing specialized equipment. This project has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No 853393-ERC-ATTIDA and from the Israel Science Foundation (ISF) under grant agreement No 1504/20. We also acknowledge the Helen Diller Quantum Center and the Russell Berrie Nanotechnology Institute at the Technion for partial financial support.
Ethics declarations
The authors declare no competing interests.